OpenStudy (anonymous):

Integrate the function. I think it would be better to change to polar coordinates.

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{-\sqrt{1-x^2}}^{\sqrt{1-x^2}} e^{-(x^2+y^2)} dydx\]

OpenStudy (anonymous):

whoa wow

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

why would you want to change it

OpenStudy (anonymous):

take the partial derivative with respects to y and then do the first integral

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\int\limits_{-\sqrt{1-x^{2}}}^{\sqrt{1-x^{2}}}e^{-(x^{2}+y^{2})}dydx=\int\limits_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int\limits_{0}^{1}e^{-r^2}rdrd \theta=\frac{(1-e^{-1})\pi}{2}\]